OpenStudy (anonymous):

Which of the following are one-to-one functions? (Select all that apply.) h(x)=IxI+4 d(x)=x^2+7 g(x)=1/(x+1) f(x)=x^3+4 Please explain your answers too.

OpenStudy (asnaseer):

one-to-one implies that 1 value in the domain of x maps to exactly one value in the domain of the function of x and vice-versa.

OpenStudy (asnaseer):

so h(x) for example is NOT a one-to-one function as h(-1) = 5 AND h(1) = 5, which means a 5 in the domain of the function does not map to exactly one point in the domain of x.

OpenStudy (amistre64):

its hard to say if 1/x is a function; but that depends on your domain of definition i spose; quads are never one 1-1, and the | | is just a quad in disguise

OpenStudy (asnaseer):

you should, hopefully, be able to us this information to answer this question now.

OpenStudy (anonymous):

Not really

OpenStudy (asnaseer):

ok - for each function, can you think of a value for x which would give you the same value for the function? e.g. h(x) gives 5 for x=-1 or x=1 any such function will NOT be a one-to-one function

OpenStudy (anonymous):

That one thats cubed cant be it either right?

OpenStudy (asnaseer):

can you think of two values for x that will give you the same answer for f(x)?

OpenStudy (anonymous):

I'm sorry I'm a little confused right now

OpenStudy (asnaseer):

can you think of two values for x for which the function:\[f(x)=x^{3}+4\]will give you the same answer?

OpenStudy (anonymous):

Wait... I think I get what you are saying... since it's an absolute value it will be 1 either way which makes you get 5.

OpenStudy (asnaseer):

yes - I think you understand now.

OpenStudy (anonymous):

Yeah I think I do... Thanks

OpenStudy (asnaseer):

so which ones do you are one-to-one functions?

OpenStudy (anonymous):

g(x)=1/(x+1) f(x)=x^3+4

OpenStudy (asnaseer):

perfecto! you are a genius!

OpenStudy (anonymous):

lol... Nah